Coulomb energy Calculate the electrostatic potential energy between a proton and a nucleus of atomic number 71 separated by 1.73 fm. Submit Answer Tries 0/15
Answer #1

electrostatic potential energy is U = k*q1*q2/r


we know that k = 9*10^9 H/m

q1 = charge on proton = 1.6*10^-19 C

q2 = charge on nucleus = 71*1.6*10^-19 C

r = 1.73*10^-15 m

then

U = k*q1*q2/r = (9*10^9*1.6*10^-19*71*1.6*10^-19)/(1.73*10^-15)


U = 9.45*10^-12 J

if it wont works add a negative sign before the answer